OpenStudy (anonymous):

The area of a regular hexagon is 30in^2. Find the length of a side. Round your answer to the nearest tenth.

OpenStudy (anonymous):

first of all u must know the formula of regular hexagon

OpenStudy (anonymous):

The area of a hexagon is given by this equation: A=2\[3\sqrt{3}/2 * L\], where L is the length of one side of the hexagon. For a hexagon, the length of one side is equal to the lengths of all the sides. For reference, go to this link: http://homepage.mac.com/terhorab/iblog/B1032916816/C181698761/E174145342/Media/hexagonArea.pdf To find the length of one side of the hexagon, solve the formula for L and put in the area.

OpenStudy (anonymous):

that is

OpenStudy (anonymous):

Typo. Area is given by the formula \[3\sqrt{3}/2 * L\]

OpenStudy (anonymous):

Wow I keep making typos. A = \[3\sqrt{3}/ 2 *L^2\]